OpenStudy (anonymous):

what is the value of y in the solution to the following system of equations? 5x-3y=-3 2x-6y=-6

OpenStudy (anonymous):

Solve for y in the first equation: -3y=-3-5x y=1+(5/3)x Now that you have this value of y in terms of x, you can replace the y term in the second equation with this value so that: 2x-6(1+(5/3)x)=-6 We can then simplify and solve for x: 2x-6+10x=-6 12x-6=-6 12x=0 x=0 Plugging this value of x into any of the two equations: 5(0)-3y=-3 -3y=-3 y=1